# There and Back Again: On the relation between Noise and Image Inversions in Diffusion Models

### v3

@@ F.1 Methodology
Next, we determine histograms for each angle, approximating the probability density function for every angle ($p_{\angle{\mathbf{x_{T}}}}$,$p_{\angle{\mathbf{x_{0}}}}$,$p_{\angle{\mathbf{\hat{x}_{T}}}}$) binned up to the precision of one degree, see fig. 11. Finally, for all angles triples candidates (where $\angle{\mathbf{x_{T}}}+\angle{\mathbf{x_{0}}}+\angle{\mathbf{\hat{x}_{T}}}=180^{\circ}$), we calculate the probability of a triangle as the product of the probabilities and choose the triplet maximizing such joint probability:
@@
F.2 Impact of T on noise-image-latent triangles
In table 12 we show how the geometrical relations between Gaussian noise, generated images and DDIM latents change when sampling and inversion is done with varying $T$.
Model $T$ $\angle{\mathbf{x_{0}}}$ $\angle{\mathbf{x_{T}}}$ $\angle{\mathbf{\hat{x}_{T}}}$
ADM $32\times 32$ $10$ $44$ $16$ $120$
$100$ $29$ $28$ $123$
$1000$ $20$ $45$ $115$
ADM $64\times 64$ $10$ $30$ $31$ $119$
$100$ $11$ $60$ $109$
$1000$ $6$ $79$ $95$
ADM $256\times 256$ $10$ $24$ $50$ $106$
$100$ $24$ $73$ $83$
$1000$ $23$ $73$ $84$
LDM $256\times 256$ $10$ $23$ $53$ $104$
$100$ $2$ $76$ $102$
$1000$ $1$ $83$ $96$
DiT $256\times 256$ $10$ $27$ $47$ $106$
$100$ $4$ $66$ $110$
$1000$ $1$ $80$ $99$
F.3 Example histograms
@@
Appendix G Plain surface thresholding
@@
Appendix H Calculating inversion error
@@
First, we generate $N=4$k images with $T=50$ diffusion steps with both ADM-64 and DiT models, saving intermediate noise predictions $\{\mathcal{E}_{t}^{S}\}_{t=50\dots 1}$ during sampling. Next, we collect diffusion model outputs during the inversion process $\{\mathcal{E}_{t}^{I}\}_{t=1\dots 50}$ assuming all the previous steps were correct. While for $t=1$, $\mathcal{E}_{1}^{I}$ can be set to the diffusion model prediction from the first inversion step $\epsilon_{\theta}(x_{0})$, for further steps, more advanced methodology is necessary. To this end, for each $t^{\prime}=2\dots 50$, we map the images to latents with DDIM Inversion, replacing in $t=1\dots t^{\prime}-1$ the predicted noise $\epsilon_{\theta}(x_{t})$ with the ground truth prediction $\mathcal{E}_{t}^{S}$ from denoising process, and collect the model output for the $t^{\prime}$ step as $\mathcal{E}_{t^{\prime}}^{I}\coloneqq\epsilon_{\theta}(x_{t^{\prime}})$. For step $t$, this methodology is equivalent to starting the inversion process from $(t+1)$ denoising step and collecting the first diffusion model prediction.
@@
Further, for the images in the dataset, we obtain binary masks $\mathcal{M}_{p}$ and $\mathcal{M}_{n}$ indicating, respectively, plain ($p$) and non-plain ($n$) pixels in the image, according to the procedure described in appendix G. To ensure that the level of noise that DM predicts in each step does not bias the results, we divide the absolute errors in each step by $l_{1}$-norm of model outputs, calculated separately for each diffusion step. For plain ($p$) pixels, it can be described as
@@
¯$\bar{\mathcal{E}}_{(t_{s},t_{e})}^{Ep}=\sum_{t^{\prime}=t_{s}}^{t_{e}}\mathcal{E}_{t}^{Ep}.$ (15)
Appendix I Pseudocode for DDIM Inversion with Forward Diffusion
@@
4:$\mathbf{\hat{x}_{f}}\leftarrow\sqrt{\bar{\alpha}_{f}}\cdot\mathbf{x_{0}}+\sqrt{1-\bar{\alpha}_{f}}\cdot\tilde{\epsilon}$triangleright forward diffusion
@@
6:     $\mathbf{\hat{x}_{t}}\leftarrow\sqrt{\alpha_{t}}\cdot\mathbf{\hat{x}_{t-1}}+(\sqrt{1-\bar{\alpha}_{t}}-\sqrt{\alpha_{t}})\cdot\epsilon_{\theta}^{(t)}(\mathbf{\hat{x}_{t-1}})$triangleright DDIM Inversion step
@@
Appendix J Details on diversity and alignment of edition
In this section, we provide results for measuring the diversity (against source images $I_{S}$) and alignment to conditioning prompts (source $P_{S}$ and target $P_{T}$) of images generated from Gaussian noise ${\mathbf{x_{T}}}$, DDIM latents ${\mathbf{\hat{x}_{T}}}$, or the latents produced by our method (described in algorithm 1) with modified prompts. We use distance-based metrics (LPIPS [61], DreamSim [11]), similarity metrics (SSIM [57], DINO [5]) to measure the diversity between source (input) images and target (edited) images, as well as the similarity between embeddings produced by the CLIP [44] model to calculate the alignment between the results and prompts. The results in for introduced latent decorrelation procedure are obtained with varying percentages of the first DDIM inversion steps replaced with forward diffusion.
In table 13, we show that by selecting a small fraction of inversion steps to replace with forward diffusion (from $2\%$ up to $6\%$), the resulting latents are more editable.
@@
by forward diff. (%$\%T$) DreamSim $\uparrow$ LPIPS $\uparrow$ SSIM $\downarrow$ DINO $\downarrow$ $P_{S}\downarrow$ $P_{T}\uparrow$
@@
by forward diff. (%$\%T$) DreamSim $\uparrow$ LPIPS $\uparrow$ SSIM $\downarrow$ DINO $\downarrow$ $P_{S}\downarrow$ $P_{T}\uparrow$
@@
Appendix K Stochastic editing of real images
@@
As preserving original image structure during editing is stated as a more difficult task for real images than the one naturally generated by diffusion model, we follow Hertz et al. [14] by, first, denoising latent encodings with source prompt (the one used during inversion), and, after $6\%$ of the steps, using target prompt as conditioning. The examples presented in fig. 13 indicate, that our approach (1) enables stochastic editing of images and (2) enables image manipulations in plain image regions, contrary to editing with naïve DDIM latents.
@@
Appendix L Noise-to-image mapping
We showcase an additional study showing the differences that occur between noise and latent encodings, from the perspective of their mapping to the images. Several works investigate interesting properties between the initial random noise and generations that result from the training objective of DDPMs and score-based models. Kadkhodaie et al. [21] show that due to inductive biases of denoising models, different DDPMs trained on similar datasets converge to almost identical solutions. This idea is further explored by Zhang et al. [59], observing that even models with different architectures converge to the same score function and, hence, the same noise-to-imape mapping. Khrulkov and Oseledets [22] show that diffusion models’ encoder map coincides with the optimal transport (OT) map when modeling simple distributions. However, other works [25, 28] contradict this finding.
L.1 Smallest $l_{2}$ mapping
Diffusion models converge to the same mapping between the Gaussian noise (${\mathbf{x_{T}}}$) and the generated images (${\mathbf{x_{0}}}$) independently on the random seed, dataset parts [21], or the model architecture [59]. We further investigate this property from the noise-sample and latent-sample mapping perspective.
In our experimental setup, we start by generating $N=2000$ images ${\mathbf{x_{0}}}$ from Gaussian noise ${\mathbf{x_{T}}}$ with a DDIM sampler and invert the images to latents ${\mathbf{\hat{x}_{T}}}$ with naïve DDIM inversion. Next, we predict resulting images for the starting noise samples (${\mathbf{x_{T}}}\rightarrow{\mathbf{x_{0}}}$) by iterating over all the $N$ noises, and for each of them, we calculate its pixel distances to all the $N$ generations. For given noise, we select the image to which such $l_{2}$-distance is the smallest. Similarly, we investigate image-to-noise (${\mathbf{x_{0}}}\rightarrow{\mathbf{x_{T}}}$), image-to-latent (${\mathbf{x_{0}}}\rightarrow{\mathbf{\hat{x}_{T}}}$) and latent-to-image (${\mathbf{\hat{x}_{T}}}\rightarrow{\mathbf{x_{0}}}$) mappings. We calculate the distance between two objects as $l_{2}$ norm of the matrix of differences between them (with $C\times H\times W$ being the dimensions of either pixel or latent space of diffusion model) as ${||x-y||}_{2}=\sqrt{\Sigma_{i}^{C}\Sigma_{j}^{H}\Sigma_{j}^{W}(x_{i,j,k}-y_{i,j,k})^{2}}$.
@@
For both noise $x^{T}$ and latents $\hat{x}^{T}$, their assignment to images in both directions can be successfully done when the denoising is performed in the latent space, as shown for DiT and LDM models. We hypothesize that this fact is connected with the KL regularization term that imposes a slight penalty towards a standard normal distribution $\mathcal{N}(0,I)$ on the learned latent during training [45].
L.2 Asymmetry of Noise-to-Image mapping
@@ Conclusion.
Those findings, connected with the experiment showing reduced diversity of latents (table 4), suggest that the DDIM latents, unlike noise, cannot be accurately assigned to samples, as the approximation error brings them towards the mean, reducing their diversity and making them closest to most of the images.
Appendix M On Noise-Image-Latent relations during diffusion training
To further explore the relationships that exist between noises, generations, and latents, we study how the relationships between them change with the training of the diffusion model. We train two diffusion models from scratch and follow the setup from Nichol and Dhariwal [39] for two unconditional ADMs for the ImageNet ($64\times 64$) and CIFAR-10 ($32\times 32$) datasets. The CIFAR-10 model is trained for $700$K steps, while the ImageNet model – for $1.5$M steps, both with a batch size of $128$. Models employ a cosine scheduler with $4$K diffusion steps.
M.1 Spatial relations of noise and latents over training time
We conclude our latent localization experiments (section 3.2) by showing that our observations are persistent across the diffusion model training process. We generate $N=2048$ images with the final model, using implicit sampling with $T=100$ steps, and invert them to the corresponding latents using checkpoints saved during the training.
In fig. 16, we show that both the angle adjacent to the noise $\angle{\mathbf{x_{T}}}$ and the distance between the latent and noise $\lVert{\mathbf{\hat{x}_{T}}}-{\mathbf{x_{T}}}\rVert_{2}$ quickly converge to the point that remains unchanged through the rest of the training, indicating that the relation between noises, latents, and samples is defined at the very early stage of the training. Additionally, we observe that the noise reconstruction error in DDIM Inversion does not degrade with the training progress.
@@
M.2 Image-to-noise distance mapping over training time
We analyze the image-noise mapping with $l_{2}$-distance from appendix L over diffusion model training time. We sample $N=2000$ Gaussian noises and generate images from them using ADM models with $T=100$ diffusion steps, calculating the accuracy of assigning images to corresponding noises (and vice versa) using the smallest $l_{2}$-distance. In fig. 17, we can observe, for both models, that the distance between noises and their corresponding generations accurately defines the assignment of initial noises given the generated samples (${\mathbf{x_{0}}}\to{\mathbf{x_{T}}}$) from the beginning of the training till the end. At the same time, the accurate reverse assignment (${\mathbf{x_{T}}}\to{\mathbf{x_{0}}}$) can only be observed at the beginning of the training when the trained model is not yet capable of generating properly formed images. Already in the beginning phase of model training, the quality of noise to image mapping rapidly drops and does not change until the end of training.
@@
M.3 Image alignment over training time
@@
Appendix N Additional results
@@
N.2 Interpolation
In section 4, we presented that interpolating DDIM latents with SLERP [48] leads to decrease in image quality and diversity when compared to Gaussian noise. In fig. 23, we qualitatively compare our method for removing correlations in latent encodings with naïve DDIM inversion in the task of image interpolation.
Figure 23: Qualitative comparison of images generated from interpolated latents produced with DDIM Inversion and our approach. Contrary to naïve DDIM inversion, proposed solution enables generating high quality objects with pixel-diverse background.

### v4

@@ H.1 Methodology
Next, we determine histograms for each angle, approximating the probability density function for every angle ($p_{\angle{\mathbf{x_{T}}}}$$p_{\angle{\mathbf{x_{0}}}}$$p_{\angle{\mathbf{\hat{x}_{T}}}}$) binned up to the precision of one degree, see Fig. 11. Finally, for all angles triples candidates (where $\angle{\mathbf{x_{T}}}+\angle{\mathbf{x_{0}}}+\angle{\mathbf{\hat{x}_{T}}}=180^{\circ}$), we calculate the probability of a triangle as the product of the probabilities and choose the triplet maximizing such joint probability:
@@
H.2 Results
Results of the experiment in Table 10 show that the angle located at the image and noise vertices (accordingly $\angle{\mathbf{x_{0}}}$ and $\angle{\mathbf{x_{T}}}$) are always acute and, in almost every case, the angle by the latent vertex ($\angle{\mathbf{\hat{x}_{T}}}$) is obtuse. This property implies that, due to approximation errors in the reverse DDIM process, latents reside in proximity to, but with a measurable offset, from the shortest-path trajectory between the noise distribution and the generated image.
Model $T$ $\angle{\mathbf{x_{0}}}$ $\angle{\mathbf{x_{T}}}$ $\angle{\mathbf{\hat{x}_{T}}}$
ADM
$32\times 32$
$10$ $44$ $16$ $120$
$100$ $29$ $28$ $123$
$1000$ 2020 $45$ $115$
ADM
$64\times 64$
$10$ $30$ $31$ $119$
$100$ $11$ $60$ $109$
$1000$ 66 $79$ $95$
ADM
$256\times 256$
$10$ $24$ $50$ $106$
$100$ $24$ $73$ $83$
$1000$ $23$ $73$ $84$
LDM
$256\times 256$
$10$ $23$ $53$ $104$
$100$ 22 $76$ $102$
$1000$ 11 $83$ $96$
DiT
$256\times 256$
$10$ $27$ $47$ $106$
$100$ 44 $66$ $110$
$1000$ 11 $80$ $99$
@@ H.2 Results
H.3 Example histograms
@@
Appendix I Plain surface thresholding
@@
Appendix J Calculating inversion error
@@
First, we generate N=4N=4k images with $T=50$ diffusion steps with both ADM-64 and DiT models, saving intermediate noise predictions $\{\mathcal{E}_{t}^{S}\}_{t=50\dots 1}$ during sampling. Next, we collect diffusion model outputs during the inversion process $\{\mathcal{E}_{t}^{I}\}_{t=1\dots 50}$ assuming all the previous steps were correct. While for $t=1$, $\mathcal{E}_{1}^{I}$ can be set to the diffusion model prediction from the first inversion step $\epsilon_{\theta}(x_{0})$, for further steps, more advanced methodology is necessary. To this end, for each $t^{\prime}=2\dots 50$, we map the images to latents with DDIM Inversion, replacing in $t=1\dots t^{\prime}-1$ the predicted noise $\epsilon_{\theta}(x_{t})$ with the ground truth prediction $\mathcal{E}_{t}^{S}$ from denoising process, and collect the model output for the $t^{\prime}$ step as $\mathcal{E}_{t^{\prime}}^{I}\coloneqq\epsilon_{\theta}(x_{t^{\prime}})$. For step $t$, this methodology is equivalent to starting the inversion process from $(t+1)$ denoising step and collecting the first diffusion model prediction.
@@
Further, for the images in the dataset, we obtain binary masks $\mathcal{M}_{p}$ and $\mathcal{M}_{n}$ indicating, respectively, plain ($p$) and non-plain ($n$) pixels in the image, according to the procedure described in Appendix I. To ensure that the level of noise that DM predicts in each step does not bias the results, we divide the absolute errors in each step by $l_{1}$-norm of model outputs, calculated separately for each diffusion step. For plain ($p$) pixels, it can be described as
@@
$\bar{\mathcal{E}}_{(t_{s},t_{e})}^{Ep}=\sum_{t^{\prime}=t_{s}}^{t_{e}}\mathcal{E}_{t}^{Ep}.$ (15)
Pixel area Diffusion steps Model
ADM-64 DiT
Plain $1,2\dots 50$ $100\%$ $15.11$ $5.67$
Non-plain $12.43$ $3.16$
Plain $1,2\dots 5$ $10\%$ $9.23$ $3.41$
Non-plain $6.80$ $1.42$
Plain $6,7\dots 50$ $90\%$ $5.87$ $2.23$
Non-plain $5.63$ $1.74$
Table 11: Total per‑pixel inversion error (normalized) over different timestep ranges $\bar{\mathcal{E}}_{(t_{s},t_{e})}^{E}$ for plain and non-plain areas. Inversion approximation error is higher for pixels related to plain image areas, especially in the first 10% of inversion steps.
Appendix K Pseudocode for DDIM Inversion with Forward Diffusion
@@
[…]
@@
6:  $\mathbf{\hat{x}_{t}}\leftarrow\sqrt{\alpha_{t}}\cdot\mathbf{\hat{x}_{t-1}}+(\sqrt{1-\bar{\alpha}_{t}}-\sqrt{\alpha_{t}})\cdot\epsilon_{\theta}^{(t)}(\mathbf{\hat{x}_{t-1}})$ $\triangleright$ DDIM Inversion step
@@
Appendix L Details on diversity and alignment of edition
In this section, we provide results for measuring the diversity (against source images $I_{S}$) and alignment to conditioning prompts (source $P_{S}$ and target $P_{T}$) of images generated from Gaussian noise $\mathbf{x_{T}}$}, DDIM latents $\mathbf{\hat{x}_{T}}$}, or the latents produced by our fix (described in Algorithm 1) with modified prompts. We use distance-based metrics (LPIPS (Zhang et al., 2018), DreamSim (Fu et al., 2023)), similarity metrics (SSIM (Wang et al., 2004), DINO (Darcet et al., 2024)) to measure the diversity between source (input) images and target (edited) images, as well as the similarity between embeddings produced by the CLIP (Radford et al., 2021) model to calculate the alignment between the results and prompts. The results in for introduced latent decorrelation procedure are obtained with varying percentages of the first DDIM inversion steps replaced with forward diffusion. In Table 12, we show that by selecting a small fraction of inversion steps to replace with forward diffusion (from $2\%$ up to $6\%$), the resulting latents are more editable.
@@
by forward diff. ($\%T$) DreamSim $\uparrow$ LPIPS $\uparrow$ SSIM $\downarrow$ DINO $\downarrow$ $P_{S}\downarrow$ $P_{T}\uparrow$
@@
by forward diff. ($\%T$) DreamSim $\uparrow$ LPIPS $\uparrow$ SSIM $\downarrow$ DINO $\downarrow$ $P_{S}\downarrow$ $P_{T}\uparrow$
@@
Appendix M Stochastic editing of real images
@@
As preserving original image structure during editing is stated as a more difficult task for real images than the one naturally generated by diffusion model, we follow Hertz et al. (2022) by, first, denoising latent encodings with source prompt (the one used during inversion), and, after $6\%$ of the steps, using target prompt as conditioning. The examples presented in Fig. 13 indicate, that our fix (1) enables stochastic editing of images and (2) enables image manipulations in plain image regions, contrary to editing with naïve DDIM latents.
@@
Appendix N Noise-to-image mapping
We showcase an additional study showing the differences that occur between noise and latent encodings, from the perspective of their mapping to the images. Several works investigate interesting properties between the initial random noise and generations that result from the training objective of DDPMs and score-based models. Kadkhodaie et al. (2024) show that due to inductive biases of denoising models, different DDPMs trained on similar datasets converge to almost identical solutions. This idea is further explored by Zhang et al. (2024a), observing that even models with different architectures converge to the same score function and, hence, the same noise-to-imape mapping. Khrulkov & Oseledets (2022) show that diffusion models’ encoder map coincides with the optimal transport (OT) map when modeling simple distributions. However, other works (Kim & Milman, 2011; Lavenant & Santambrogio, 2022) contradict this finding.
N.1 Smallest $l_{2}$ mapping
Diffusion models converge to the same mapping between the Gaussian noise ($\mathbf{x_{T}}$}) and the generated images ($\mathbf{x_{0}}$}) independently on the random seed, dataset parts (Kadkhodaie et al., 2024), or the model architecture (Zhang et al., 2024a). We further investigate this property from the noise-sample and latent-sample mapping perspective.
[…]
@@
For both noise $x^{T}$ and latents $\hat{x}^{T}$, their assignment to images in both directions can be successfully done when the denoising is performed in the latent space, as shown for DiT and LDM models. We hypothesize that this fact is connected with the KL regularization term that imposes a slight penalty towards a standard normal distribution $\mathcal{N}(0,I)$ on the latent during training (Rombach et al., 2021).
N.2 Asymmetry of Noise-to-Image mapping
@@ Conclusion.
Those findings, connected with the reduced diversity of latents (Table 3), suggest that the DDIM latents, unlike noise, cannot be accurately assigned to samples, as the error brings them towards the mean, reducing their diversity and making them closest to most of the images.
Appendix O On Noise-Image-Latent relations during diffusion training
To further explore the relationships that exist between noises, generations, and latents, we study how the relationships between them change with the training of the diffusion model. We train two diffusion models from scratch and follow the setup from Nichol & Dhariwal (2021) for two unconditional ADMs for the ImageNet ($64\times 64$) and CIFAR-10 ($32\times 32$) datasets. The CIFAR-10 model is trained for 700700K steps, while the ImageNet model – for 1.51.5M steps, both with a batch size of $128$. Models employ a cosine scheduler with 44K diffusion steps.
O.1 Spatial relations of noise and latents over training time
We conclude our latent localization experiments (Section 3.2) by showing that our observations are persistent across the diffusion model training process. We generate $N=2048$ images with the final model, using implicit sampling with $T=100$ steps, and invert them to the corresponding latents using checkpoints saved during the training. In Fig. 16, we show that both the angle adjacent to the noise $\angle{\mathbf{x_{T}}}$ and the distance between the latent and noise $\lVert{\mathbf{\hat{x}_{T}}}-{\mathbf{x_{T}}}\rVert_{2}$ quickly converge to the point that remains unchanged through the rest of the training, indicating that the relation between noises, latents, and samples is defined at the very early stage of the training. Additionally, we observe that the noise reconstruction error in DDIM Inversion does not degrade with the training progress.
@@
O.2 Image-to-noise distance mapping over training time
We analyze the image-noise mapping with $l_{2}$-distance from Appendix N over diffusion model training time. We sample $N=2000$ Gaussian noises and generate images from them using ADM models with $T=100$ diffusion steps, calculating the accuracy of assigning images to corresponding noises (and vice versa) using the smallest $l_{2}$-distance. In Fig. 17, we can observe, for both models, that the distance between noises and their corresponding generations accurately defines the assignment of initial noises given the generated samples (${\mathbf{x_{0}}}\to{\mathbf{x_{T}}}$) from the beginning of the training till the end. At the same time, the accurate reverse assignment (${\mathbf{x_{T}}}\to{\mathbf{x_{0}}}$) can only be observed at the beginning of the training when the trained model is not yet capable of generating properly formed images. Already in the beginning phase of model training, the quality of noise to image mapping rapidly drops and does not change until the end of training.
@@
O.3 Image alignment over training time
@@
Appendix P Additional results
@@
P.2 Interpolation
In Section 4, we presented that interpolating DDIM latents with SLERP (Shoemake, 1985) leads to decrease in image quality and diversity when compared to Gaussian noise. In Fig. 23, we qualitatively compare our fix for removing correlations in latent encodings with naïve DDIM inversion in the task of image interpolation.
Figure 23: Qualitative comparison of images generated from interpolated latents produced with DDIM Inversion and our fix. Contrary to naïve DDIM inversion, proposed solution enables generating high quality objects with pixel-diverse background.